It is know that the population mean for the Full Scale IQ of the WAIS is 100 with a standard deviation of 15. A researcher assesses a sample of 200 adults and find that they have a mean Full Scale IQ of 102. The point of estimate of the mean for this group is _________ and the 99% confidence interval for this group is_________.
A. 100; (97.27; 102.73)
B. 102 ;(97.27 ; 102.73)
C. 100 ; (99.27 ; 104.73)
D. 102; (99.27 ; 104.73)
It is know that the population mean for the Full Scale IQ of the WAIS is 100 with a standard deviation of 15. A researcher assesses a sample of 200 adults and find that they have a mean Full Scale IQ of 102. The point of estimate of the mean for this group is _________ and the 90% confidence interval for this group is_________.
A. 100 ; (100.26 ; 103.74)
B. 102 ; (100.26 ; 103.74)
C. 100 ; ( 98.26; 101.74)
D. 102 ; (98.26 ; 101.74)



Answer by ewatrrr(24785)


Hi
ME = 2.576*15/sqrt(200) = 2.73
CI: 102-2.73 < u < 102 + 2.73
B. 102 ;(97.27 ; 102.73)
ME = 1.645*15/sqrt(200) = 1.74
CI: 102-1.74 < u < 102 + 1.74
B. 102 ; (100.26 ; 103.74)
CI z
90% z =1.645
92% z = 1.751
95% z = 1.96
98% z = 2.326
99% z = 2.576